OpenStudy (anonymous):

If the hypotenuse of a right isosceles triangle is 20 units long, what is the length of each leg? Answer A. 10 B. 10√ 2 C. 10√ 3 D. Not enough information to find the value.

OpenStudy (anonymous):

20^2=2a^2 20=4a a=5 ength of each leg? is 5

OpenStudy (anonymous):

so is it enough information?

OpenStudy (anonymous):

sorry use pythogarus formula 20^2=a^2+a^2 400=2.a^2 200=a^2 a=root(200) a=10root(2)

OpenStudy (anonymous):

option b

OpenStudy (anonymous):

ok thanks sails :)

OpenStudy (anonymous):

giv me medal yar

OpenStudy (anonymous):

how

OpenStudy (anonymous):

click blu color best response button

OpenStudy (anonymous):

ok cool